Main Company has a fiscal year ending on December 31. The company purchased equipment costing S88.000 on October 2, 2011. The equipment was purchased by paying 30% down and signing a 9%, 120-day note payable for the balance. 1. Prepare the journal entries to record the following events. Use a 360-day year and round all amounts to the nearest dollar. The purchase of the equipment on October 2, 2011 The accrual of interest on December 31, 2011 Payment of the note on January 30, 2012 2. Determine the balance of any current liabilities associated with the note on December 31,2011
Answer #1

1.Purchase of Equipment

Date

Accounts

Debit

Credit

Oct 2 2011

Equipment

$88,000

Cash

26,400

Short term Note Payable

61600

Cash = 88000*30% = 26,400

Notes Payable = 88,000 – 26400

2. Accrual of Interest Expense

Date

Accounts

Debit

Credit

December 31

Interest Expense

1,386

Interest Payable

1,386

But the actual amount interest is needed to be given for 90 days from Oct 2 20111 to January 30 2012.

For 90 days, the amount of interest = 61600*9% 90/120 = 1386

3.Payment of Note

Date

Accounts

Debit

Credit

January 30,2012

Short term Note Payable

61600

Interest Payable

1,386

Interest Expense

462

Cash

63,448

Interest = 61,600*9%*120/360 = 1,848

Interest Expense = 1848 – 1386

Cash = Short term note payable + Interest Payable + Interest Expense

= 61600+1386+462 = 63,448

Balance of Current Liability

Note Payable

61600

Interest Payable

1386

$ 62986